# Optimizing Optimizers for Fast Gradient-Based Learning

### v1

@@ 4 Empirical demonstration
Method Test acc. % Train loss
Best baseline 52.57 $\pm$ 1.10 0.2080 $\pm$ 0.0004
Ours 52.77 $\pm$ 0.93 0.2084 $\pm$ 0.0003
Method Test acc. % Train loss
Best baseline 76.20 $\pm$ 0.33 0.1927 $\pm$ 0.0005
Ours 76.30 $\pm$ 0.31 0.1925 $\pm$ 0.0005
@@
One of the practical advantages of our framework is the automatic determination of optimal hyperparameters, as shown in Corollaries 5 and 6. The theory clarifies that optimal hyperparameters depend on the gradient distribution, as illustrated in Figure 1. Here, we provide simple instantiations of these theoretical frameworks and demonstrate their validity and practical usefulness. The implementation of the argmax in equations 12 and 16 can vary. However, the key idea is to maximize the cosine similarity among available options. The simplest way is to maintain two fixed optimizers with different hyperparameters $\beta^{(1)}$ and $\beta^{(2)}$, compute the argmax operands for each option, and select the optimizer with the largest value dynamically to obtain the parameter update.
@@
Using this simple instantiation, we empirically demonstrate the theory by training a ResNet-18 (he2016deep) model on the CIFAR-100 dataset (krizhevsky2009learning). We provide the full algorithm in Appendix B. Baseline optimizers are trained with fixed hyperparameters, following typical machine learning practices. For the momentum of SGD+Momentum, we tried $\beta\in[0.01,0.999]$ and reported the best one. For the Adam, we tried $\beta_{1}\in[0.1,0.99]$ while keeping $\beta_{2}=0.999$ fixed, and reported the best one. Figure 3 and Tables 2 and 2 summarize the results. Our automatic hyperparameter tuning shows comparable and often better performance than the baseline optimizers with fixed hyperparameters in both final accuracy and convergence speed. This demonstrates that tedious manual hyperparameter tuning is unnecessary and can be replaced by our framework.
@@
We also demonstrate our framework in more practical scenarios training large language models (LLMs) and vision transformers (ViTs). We train a Gemma-2B (gemma2024open) and Llama-3-8B (llama32024herd) using low rank adaptation (LoRA) (hu2022lora) with standard settings on the MetaMathQA-395K dataset (yu2024metamath) and compare the results with baseline optimizers of fixed hyperparameters on GSM8K (cobbe2021gsm8k). Table 4 and 4 summarize the results of ten runs for each model. Furthermore, we also train a Gemma-2B with both LoRA and full fine-tuning on the Commonsense-170K dataset (hu2023llmadapters) using ours and baseline optimizers of fixed hyperparameters. We then compare the results on various reasoning tasks such as BoolQ (clark2019boolq), PIQA (bisk2020piqa), Social IQA (sap2019socialiqa), HellaSwag (zellers2019hellaswag), WinoGrande (sakaguchi2019winogrande), and OBQA (hu2023llmadapters), and summarize the results in Table 5. Finally, we train ViT-B and ViT-L models (dosovitskiy2021image) by LoRA on various image classification tasks including Cars (krause20133d), CIFAR-100 (krizhevsky2009learning), CUB-200 (wah2011cub200), DTD (cimpoi2014describing), Food-101 (bossard2014food101), RESISC45 (cheng2017resisc45), and SUN397 (xiao2010sun). Table 6 summarizes the results. Detailed settings and extended results are provided in Appendix C.
@@
Our automatic hyperparameter tuning, again, shows comparable and often better performance than the baseline optimizers with fixed hyperparameters. This improvement is consistent across various architectures and tasks. We emphasize that the overhead of our method is only less than 5% of the training time, and even less than what is measurable for LLM training, where we encountered speed improvement instead of slowdown. Given the huge amount of time required for hyperparameter tuning, this additional cost is acceptable. In summary, our theory enables us to greatly reduce the workload of manual hyperparameter tuning in practical scenarios, where computations are scarce resources. In addition to this, we notice that our automatic framework can be combined with validation-aware tuning of optimizing environments, which is typically done by human engineers by manual inspection of validation curves. We provide a proof-of-concept in Appendix D.
5 Convergence endpoint of greedy optimal optimizers
@@
Despite the greedy objective for instantaneous progress of learning, the previous experiments reveal that the resulting solutions often yield better test accuracy as well. This section provides a theoretical foundation for this empirical observation. Note that our primary goal here is to enhance the usability of existing families of optimizers that are already verified both theoretically and empirically. That is, we assume that the optimizers under study has a well-established convergence analysis in its general form, i.e., regardless of the choice of hyperparameters. For readers who are interested in these aspects, we refer to works of ghadimi2013stochastic; yang2016unified; reddi2018convergence; zhou2018convergence; assran2020convergence; li2020high; cutkosky2020momentum; defossez2022a. We instead focus on the dependence of the learning dynamics and its target endpoint on the greedy selection of optimizer from the predefined family of optimizers.
@@
Proposition 7 (Convergence endpoint of commutative optimizers for least squares).
@@
If the loss function is not a sum of squares, e.g., a nonparametric loss function $\ell(f,y)$ with gradient $s=\nabla_{\theta}\ell(f,y)$, then $R=\mathbb{E}[gg^{\top}]=\mathbb{E}[J^{\top}ss^{\top}J]\eqqcolon J^{\top}\Sigma_{s}J$ is the gradient moment, where $\Sigma_{s}\coloneqq\mathbb{E}[ss^{\top}]$ is the gradient moment of the criterion $\ell$. Even in this case, Lemma 8 shows that $Q^{\star}$ does commute with $R$. However, $Q^{\star}$ may not commute with the Gram matrix $J^{\top}J$ in general, which prevents achieving the canonical pseudoinverse solution of Theorem 7. Nevertheless, by carefully choosing the desired families of optimizers, we can also achieve this commutativity condition $Q,J^{\top}J]=0$. Notable examples of such commuting optimizers are the families of simple gradient descent (with or without momentum) that are scalar multiples of the identity matrix, and the families of Adam optimizers sharing the same $\beta_{2}$ value, which fixes the diagonal structure of the optimizer. These are the families that used in the demonstrations in Section 3.
In short, our greedy optimal optimizer does minimal disruption to the convergence endpoint for the least squares problem by maintaining the commutativity property of Lemma 8. This effect can be seen in Figure 2(a,d), where the greedy optimal optimizer achieves the fastest convergence without suffering from the nonisotropic curvature of the problem, leading to precise minimization.
Extending this to more general problems requires a concept of kernels. Define the optimizer-augmented kernel (OAK) $K_{Q}$ induced by the optimizer $Q$ (jacot2018neural; geifman2024controlling):
@@
with time $t$. This is analogous to the parameter motion formula in Proposition 7, but now in the function space. The solution to this ODE at training points $X$ is:
@@
Theorem 9 (Convergence endpoint of greedy optimal optimizers).
[proof] Let $Q^{\star}$ solve $\max_{Q\in\mathcal{Q}}\langle Q,R\rangle$ for a Frobenius, spectral, data-metric, or diagonal families, where $R=\mathbb{E}[gg^{\top}]$. Then, the convergence endpoint $f^{\infty}_{Q^{\star}}$ is, under squared loss and OAK dynamics with small initialization, the unique minimum-norm interpolant in the RKHS $\mathcal{H}_{K_{Q^{\star}}}$:
$f^{\infty}_{Q^{\star}}(\cdot)\,=\,\arg\underset{f(X)=y}{\operatorname{min}}\ \|f\|_{\mathcal{H}_{K_{Q^{\star}}}}^{2}\!=\,K_{Q^{\star}}(\cdot,X)K_{Q^{\star}}^{\dagger}y,\quad\;\|f^{\infty}_{Q^{\star}}\|_{\mathcal{H}_{K_{Q^{\star}}}}^{2}\!=\,y^{\top}K_{Q^{\star}}^{\dagger}y\$(23)
where $K_{Q^{\star}}^{\dagger}$ is the Moore-Penrose pseudoinverse of $K_{Q^{\star}}$. In particular, whenever interpolation is possible, all such $Q$ achieve the same minimal training loss in this OAK dynamics.
6 Master table of optimizers
Throughout this work, we have derived various types of optimizers from our convex optimization framework. We can now register various optimizers under a single unified table as shown in Table 7. For detailed derivation of each optimizer family, click the first column items of the table to jump to the corresponding section. For SGD with momentum, Adam (kingma2015adam), and AdamW (loshchilov2017decoupled), see Corollaries 5 and 6 in the main text. For other optimizers, we have provided the detailed derivation in Appendix E.
7 Scope and limitations
Long-horizon objective from greedy paradigm.
In order to simplify the analysis, this work resorts to the greedy paradigm, primarily focusing on instantaneous progress of learning. As a trade-off, global optimality guarantee requires further investigation under this greedy paradigm. We have provided a theoretical analysis on the convergence targets near the linearized dynamics regime in Section 5, and rely on more sophisticated guarantees on the families of optimizers from optimizer-specific prior works. It is, therefore, recommended to combine our theoretical results with the optimizer-specific guarantees in order to use our theory to devise a new type of optimizers.
Choice of trust region types.
Rather than determining which optimizer class is optimal for a given task, this work provides an optimization framework within user-defined feasible sets. Our framework helps engineers by reducing hyperparameter search effort; however, it still requires intelligent choice of trust region types, i.e., which optimizer class fits the task.
Renovating existing optimizers.
In Theorems 1 and 3, we provide general construction of optimal optimizers from convex constraints. However, in this work, we focus on the well-established optimizers, and supplement them with a systematic methodology to find the right hyperparameters for a given task. Designing a new class of optimizers will be a natural extension of this work.
@@ 8 Conclusion
We established a firm theoretical grounding for systematically achieving optimal optimizers in a greedy sense. Our convex optimization framework connects commonly used optimizers to convex constraint sets, merging those independently developed techniques into a single unified framework in Table 7. Our main results, Theorems 1 and 3, and Lemma 4 are general tools that can be extended to arbitrary trust region to invent new families of optimizers for specific uses. Our theory, therefore, does not disprove the no free lunch theorem; rather, it provides a principled way to leverage this wisdom to flexibly design and adapt optimizers for our own problems at hand.
Ethics statement
We acknowledge the ICLR Code of Ethics.
Reproducibility statement
All the proofs of the theoretical part of this paper, including every Lemma, Theorem, Proposition, and Corollary, are provided in Appendix G with detailed derivations, starting from the basic definitions and assumptions made in the main text. Moreover, omitted theoretical results are elaborated in Appendix F. Regarding the implementation, Appendix B gives the algorithm to realize our theoretically justified optimal optimizers.
LLM usage statement
We deeply acknowledge the usefulness of LLMs in revising the manuscript, especially for fixing vocabulary and grammar-related issues. We also used LLMs to check the correctness and coherence of the proofs and notations. This greatly helped us in identifying awkward mistakes we had been making all along.
@@ Appendix A Related work
The closest work to ours is AdaReg (gupta2017unified), which presents a minimization framework for selecting the best optimizer adaptively. Specifically, AdaReg generalizes AdaGrad (duchi2011adaptive) and Online Newton Step (ONS) (hazan2007logarithmic) into solutions of a single convex optimization problem that resembles our problem P2 for the stateless case. However, this is done with a regularization term $\Phi(Q)$ that penalizes the complexity of the optimizer. Our framework extends this idea to significantly broader family of optimizers, including stateless and dynamic optimizers, with respect to general families of trust regions. This allows us to unify many existing optimizers in practice and find their optimal hyperparameters. On the other hand, Frank-Wolfe methods (frank1956algorithm; garber2021frank) considers finding the most aligned optimization step for the constrained convex optimization problem. This alignment principle is similar to our analysis in Section 5, where we generalize this idea to the practical algorithms for deep learning.
@@
Another line of work is the performance estimation problem (PEP) framework (drori2014performance; kim2017optimized_gradient; taylor2017smooth_interpolation; goujaud2024pepit_springer), where first order methods for convex optimization are categorized, compared, and suggested based on their worst-case performance. Although we share a general philosophy to algorithmically suggest the best optimizer for each task, our greedy paradigm is orthogonal to the PEP framework, as we focus on the instantaneous performance of the optimizers in general gradient-based learning. We also encompass a broader family of optimizers, unifying existing widely-used optimizers in deep learning such as SGD with Nesterov momentum (nesterov1983method), AdamW (loshchilov2017decoupled), LAMB (you2020large), K-FAC (martens2015optimizing), Shampoo (gupta2018shampoo), and Lion (chen2023symbolic).
@@
Techniques like symbolic discovery (chen2023symbolic; zheng2022symbolic), non-parametric optimizer search (wang2022efficient), and neural optimizer search (bello2017neural) are also related to our work, as their objective is to discover the optimal optimizer for a given task. In their framework, symbolic optimizers are obtained by a tree-based search of a predefined set of optimizers. Ours instead lets the engineer select the broader family of optimizers, and then provides a mathematical tool to find the optimal solution among them. Therefore, these works are also orthogonal to ours.
@@
Many works have proposed to automatically tune the hyperparameters governing optimization. Most of them adopt a learning framework to find a good set of hyperparameters including learning rates (daniel2016stepsize), their schedules (xu2017rl; xu2019adaptive), and other optimizer parameters (shaban2019truncated). Hypergradient methods (maclaurin2015reversible; baydin2017hypergradient; grazzi2020hypergradient; moskovitz2019fop) are also proposed to find the optimal hyperparameters. Instead of resorting to learning-based methods, we establish a theoretical framework through the lens of convex optimization problems (boyd2004convex). By doing so, we can classify well-used optimizers such as SGD with momentum and Adam (kingma2015adam) as special cases of our framework, and provide a systematic way to determine the optimal hyperparameters for these optimizers.
@@
Learning to optimize (li2016learning) aims to adapt the optimizer to a given task by treating optimizers as learnable parametric models (andrychowicz2016learning). Various architectures have been explored, including RNNs (andrychowicz2016learning; wichrowska2017learned; lv2017learning), Transformers (chen2022towards; moudgil2023learning; jain2024mnemosyne), and per-tensor HyperNetworks (ha2016hypernetworks; metz2022velo). Their primary focus is on meta-training these optimizer-networks for stability and adaptability. These works represent a nontraditional, network-based family of generally nonconvex optimizers, which is not generally compatible with our framework which is based on convex optimization.
@@
Rooted in the human-inspired philosophy (schmidhuber1987srl; bengio1990synaptic), meta-learning is another line of work that shares a similar spirit with learning to optimize (gharoun2023meta). A large proportion of works on meta-learning target few-shot learning tasks, which prepare the model, not the optimizer, for downstream tasks (vinyals2016matching; finn2017model; yu2024survey; sun2019meta). Among them, Meta-SGD (li2017meta) is noteworthy, as it prepares the optimizer. However, the problem set we address is general gradient-based learning, which differs from the tasks of concern in meta-learning.

### v2

@@ 4 Empirical Demonstration
Method Test acc. % Train loss
Best baseline 52.57 $\pm$ 1.10 0.2080 $\pm$ 0.0004
Ours 52.77 $\pm$ 0.93 0.2084 $\pm$ 0.0003
Method Test acc. % Train loss
Best baseline 76.20 $\pm$ 0.33 0.1927 $\pm$ 0.0005
Ours 76.30 $\pm$ 0.31 0.1925 $\pm$ 0.0005
@@
One of the practical advantages of our framework is the automatic determination of optimal hyperparameters, as shown in Corollaries 3.3 and 3.4. The theory clarifies that optimal hyperparameters depend on the gradient distribution. Here, we provide simple instantiations of these theoretical frameworks and demonstrate their validity and practical usefulness. The implementation of the argmax in equations 13 and 18 can vary. However, the key idea is to maximize the cosine similarity among available options. The simplest way is to maintain two fixed optimizers with different hyperparameters $\beta^{(1)}$ and $\beta^{(2)}$, compute the argmax operands for each option, and select the optimizer with the largest value dynamically to obtain the parameter update.
@@
Using this simple instantiation, we empirically demonstrate the theory by training a ResNet-18 (He et al., 2016) model on the CIFAR-100 dataset (Krizhevsky, 2009). We provide the full algorithm in Appendix B. Baseline optimizers are trained with fixed hyperparameters, following typical machine learning practices. For the momentum of SGD+Momentum, we tried $\beta\in[0.01,0.999]$ and reported the best one. For the Adam, we tried $\beta_{1}\in[0.1,0.99]$ while keeping $\beta_{2}=0.999$ fixed, and reported the best one. Figure 3 and Tables 2 and 2 summarize the results. Our automatic hyperparameter tuning shows comparable and often better performance than the baseline optimizers with fixed hyperparameters in both final accuracy and convergence speed. This demonstrates that tedious manual hyperparameter tuning is unnecessary and can be replaced by our framework.
@@
We also demonstrate our framework in more practical scenarios training large language models (LLMs) and vision transformers (ViTs). We train a Gemma-2B (Gemma Team and others, 2023) and Llama-3-8B (Grattafiori and others, 2024) using low rank adaptation (LoRA) (Hu et al., 2022) with standard settings on the MetaMathQA-395K dataset (Yu et al., 2024a) and compare the results with baseline optimizers of fixed hyperparameters on GSM8K (Cobbe et al., 2021). Table 6 and 6 summarize the results of ten runs for each model. Furthermore, we also train a Gemma-2B with both LoRA and full fine-tuning on the Commonsense-170K dataset (Hu et al., 2023) using ours and baseline optimizers of fixed hyperparameters. We then compare the results on various reasoning tasks such as BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Social IQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), and OBQA (Hu et al., 2023), and summarize the results in Table 3. Finally, we train ViT-B and ViT-L models (Dosovitskiy et al., 2021) by LoRA on various image classification tasks including Cars (Krause et al., 2013), CIFAR-100 (Krizhevsky, 2009), CUB-200 (Wah et al., 2011), DTD (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), RESISC45 (Cheng et al., 2017), and SUN397 (Xiao et al., 2010). Table 4 summarizes the results. Detailed settings and extended results are provided in Appendix C.
@@
Our automatic hyperparameter tuning, again, shows comparable and often better performance than the baseline optimizers with fixed hyperparameters. This improvement is consistent across various architectures and tasks. We emphasize that the overhead of our method is only less than 5% of the training time, and even less than what is measurable for LLM training. Given the huge amount of time required for hyperparameter tuning, this additional cost is acceptable. In summary, our theory enables us to greatly reduce the workload of manual hyperparameter tuning in practical scenarios, where computations are scarce resources. In addition to this, we notice that our automatic framework can be combined with validation-aware tuning of optimizing environments, which is typically done by human engineers by manual inspection of validation curves. We provide a proof-of-concept in Appendix D.
5 Convergence Endpoint Analysis
@@
Despite the greedy objective for instantaneous progress of learning, the previous experiments reveal that the resulting solutions often yield better test accuracy as well. This section provides a theoretical foundation for this empirical observation. Note that our primary goal here is to enhance the usability of existing families of optimizers that are already verified both theoretically and empirically. That is, we assume that the optimizers under study has a well-established convergence analysis in its general form, i.e., regardless of the choice of hyperparameters. For readers who are interested in these aspects, we refer to works of Ghadimi and Lan (2013); Yang et al. (2016); Reddi et al. (2018); Zhou et al. (2024); Assran and Rabbat (2020); Li and Orabona (2020); Cutkosky and Mehta (2020); Defóssez et al. (2022). We instead focus on the dependence of the learning dynamics and its target endpoint on the greedy selection of optimizer from the predefined family of optimizers.
@@
Proposition 5.1 (Convergence endpoint of commutative optimizers for least squares).
@@
If the loss function is not a sum of squares, e.g., a nonparametric loss function $\ell(f,y)$ with gradient $s=\nabla_{\theta}\ell(f,y)$, then $R=\mathbb{E}[gg^{\top}]=\mathbb{E}[J^{\top}ss^{\top}J]\eqqcolon J^{\top}\Sigma_{s}J$ is the gradient moment, where $\Sigma_{s}\coloneqq\mathbb{E}[ss^{\top}]$ is the gradient moment of the criterion $\ell$. Even in this case, Lemma 5.2 shows that $Q^{\star}$ does commute with $R$. However, $Q^{\star}$ may not commute with the Gram matrix $J^{\top}J$ in general, which prevents achieving the canonical pseudoinverse solution of Theorem 5.1. Nevertheless, by carefully choosing the desired families of optimizers, we can also achieve this commutativity condition $Q,J^{\top}J]=0$. Notable examples of such commuting optimizers are the families of simple gradient descent (with or without momentum) that are scalar multiples of the identity matrix, and the families of Adam optimizers sharing the same $\beta_{2}$ value, which fixes the diagonal structure of the optimizer. These are the families that used in the demonstrations in Section 3.
In short, our greedy optimal optimizer does minimal disruption to the convergence endpoint for the least squares problem by maintaining the commutativity property of Lemma 5.2. This effect can be seen in Figure 2(a,d), where the greedy optimal optimizer achieves the fastest convergence without suffering from the nonisotropic curvature of the problem, leading to precise minimization.
Extending this to more general problems requires a concept of kernels. Define the optimizer-augmented kernel (OAK) $K_{Q}$ induced by the optimizer $Q$ (Jacot et al., 2018; Geifman et al., 2024):
@@
with time $t$. This is analogous to the parameter motion formula in Proposition 5.1, but now in the function space. The solution to this ODE at training points $X$ is:
@@
Theorem 5.3 (Convergence endpoint of greedy optimal optimizers).
[proof] Let $Q^{\star}$ solve $\max_{Q\in\mathcal{Q}}\langle Q,R\rangle$ for Frobenius, spectral, data-metric, or diagonal families, where $R=\mathbb{E}[gg^{\top}]$. Then, the convergence endpoint $f^{\infty}_{Q^{\star}}$ is, under squared loss and OAK dynamics with small initialization, the unique minimum-norm interpolant in the RKHS $\mathcal{H}_{K_{Q^{\star}}}$:
$f^{\infty}_{Q^{\star}}(\cdot)=\arg\underset{f(X)=y}{\operatorname{min}}\ \|f\|_{\mathcal{H}_{K_{Q^{\star}}}}^{2}\!\!=K_{Q^{\star}}(\cdot,X)K_{Q^{\star}}^{\dagger}y,$ (25)
$\|f^{\infty}_{Q^{\star}}\|_{\mathcal{H}_{K_{Q^{\star}}}}^{2}\!=\,y^{\top}K_{Q^{\star}}^{\dagger}y$ (26)
where $K_{Q^{\star}}^{\dagger}$ is the Moore-Penrose pseudoinverse of $K_{Q^{\star}}$. All such $Q$ achieve the same minimal training loss in this OAK dynamics, whenever interpolation is possible.
6 Master Table of Optimizers
@@
Table 8: Simple analogy between classical mechanics and gradient-based learning. The optimizer plays the role of a constitutive law, mediating how the “force” (gradient) induces “motion” (updates) on the system.
Physics Gradient-Based Learning Meaning
Position $x$ Parameters $\theta$ What moves / changes
Velocity $v=\dot{x}$ Parameter update $u=\dot{\theta}$ How it moves / deforms
Energy / Potential $U$ Loss $\mathcal{L}$} Objective to minimize
Force $f=-\nabla_{x}U$ Gradient $g=-\nabla_{\theta}{\mathcal{L}}$ Driving signal
Constitutive law $f=M(t,x,\dot{x},\ddot{x})$ Optimizer $Q$ Converts force to motion
[…]
Throughout this work, we have derived various types of optimizers from our convex optimization framework and the physics analogy described in Section 1. This analogy is also summarized in Table 8. In our first-order optimization scenario, where gradients are linearly transformed into parameter updates, the constitutive law $Q$ is similar to the governing equation of damped motion in a viscous fluid, f=bvf=-bv, but with the sign of the factor $b$ inverted. This analogy is quite general, and can be extended to various types of physics, including those involving acceleration and even higher-order dynamics. However, this is beyond our current scope, and we hope to explore these possibilities in future work.
We can now organize various optimizers in a single unified table, as shown in Table 7. For detailed derivations of each optimizer family, click the items in the first column of the table to jump to the corresponding section. For SGD with momentum, Adam (Kingma and Ba, 2015), and AdamW (Loshchilov and Hutter, 2019), see Corollaries 3.3 and 3.4 in the main text. For other optimizers, we have provided the corresponding theorems and detailed derivations in Appendix E.
@@ 7 Conclusion
We established a firm theoretical grounding for systematically achieving optimal optimizers in a greedy sense. Our convex optimization framework connects commonly used optimizers to convex constraint sets, merging those independently developed techniques into a single unified framework in Table 7. Our main results, Theorems 2.1 and 3.1, and Lemma 3.2 are general tools that can be extended to arbitrary trust region to invent new families of optimizers for specific uses. Our theory, therefore, does not disprove the no free lunch theorem; rather, it provides a principled way to leverage this wisdom to flexibly design and adapt optimizers for our own problems at hand.
Software and Data
The code of this paper is available at https://github.com/ironjr/loot.
Impact Statement
The main goal of this paper is two-fold: (1) to reduce the wasteful time and budget currently spent on hyperparameter tuning, and (2) to broaden the accessibility of machine learning techniques by alleviating the need for such efforts. By advancing this direction, we expect to lower the computational overhead and operational costs of large-scale machine learning systems. Ultimately, we hope this contributes to a more democratic and sustainable AI ecosystem, benefiting both society and the environment.
@@ Appendix A Related Work
The closest work to ours is AdaReg (Gupta et al., 2017), which presents a minimization framework for selecting the best optimizer adaptively. Specifically, AdaReg generalizes AdaGrad (Duchi et al., 2011) and Online Newton Step (ONS) (Hazan et al., 2007) into solutions of a single convex optimization problem that resembles our problem P2 for the stateless case. However, this is done with a regularization term $\Phi(Q)$ that penalizes the complexity of the optimizer. Our framework extends this idea to significantly broader family of optimizers, including stateless and dynamic optimizers, with respect to general families of trust regions. This allows us to unify many existing optimizers in practice and find their optimal hyperparameters. On the other hand, Frank-Wolfe methods (Frank and Wolfe, 1956; Garber and Wolf, 2021) considers finding the most aligned optimization step for the constrained convex optimization problem. This alignment principle is similar to our analysis in Section 5, where we generalize this idea to the practical algorithms for deep learning.
@@
Another line of work is the performance estimation problem (PEP) framework (Drori and Teboulle, 2014; Kim and Fessler, 2017; Taylor et al., 2017; Goujaud et al., 2024), where first order methods for convex optimization are categorized, compared, and suggested based on their worst-case performance. Although we share a general philosophy to algorithmically suggest the best optimizer for each task, our greedy paradigm is orthogonal to the PEP framework, as we focus on the instantaneous performance of the optimizers in general gradient-based learning. We also encompass a broader family of optimizers, unifying existing widely-used optimizers in deep learning such as SGD with Nesterov momentum (Nesterov, 1983), AdamW (Loshchilov and Hutter, 2019), LAMB (You et al., 2020), K-FAC (Martens and Grosse, 2015), Shampoo (Gupta et al., 2018), and Lion (Chen et al., 2023).
@@
Techniques like symbolic discovery (Chen et al., 2023; Zheng et al., 2022), non-parametric optimizer search (Wang et al., 2022), and neural optimizer search (Bello et al., 2017) are also related to our work, as their objective is to discover the optimal optimizer for a given task. In their framework, symbolic optimizers are obtained by a tree-based search of a predefined set of optimizers. Ours instead lets the engineer select the broader family of optimizers, and then provides a mathematical tool to find the optimal solution among them. Therefore, these works are also orthogonal to ours.
@@
Many works have proposed to automatically tune the hyperparameters governing optimization. Most of them adopt a learning framework to find a good set of hyperparameters including learning rates (Daniel et al., 2016), their schedules (Xu et al., 2017, 2019), and other optimizer parameters (Shaban et al., 2019). Hypergradient methods (Maclaurin et al., 2015; Baydin et al., 2017; Grazzi et al., 2020; Moskovitz et al., 2019) are also proposed to find the optimal hyperparameters. Instead of resorting to learning-based methods, we establish a theoretical framework through the lens of convex optimization problems (Boyd and Vandenberghe, 2004). By doing so, we can classify well-used optimizers such as SGD with momentum and Adam (Kingma and Ba, 2015) as special cases of our framework, and provide a systematic way to determine the optimal hyperparameters for these optimizers.
@@
Learning to optimize (Li and Malik, 2016) aims to adapt the optimizer to a given task by treating optimizers as learnable parametric models (Andrychowicz et al., 2016). Various architectures have been explored, including RNNs (Andrychowicz et al., 2016; Wichrowska et al., 2017; Lv et al., 2017), Transformers (Chen et al., 2022; Moudgil et al., 2023; Jain et al., 2024), and per-tensor HyperNetworks (Ha et al., 2016; Metz et al., 2022). Their primary focus is on meta-training these optimizer-networks for stability and adaptability. These works represent a nontraditional, network-based family of generally nonconvex optimizers, which is not generally compatible with our framework which is based on convex optimization.
@@
Rooted in the human-inspired philosophy (Schmidhuber, 1987; Bengio et al., 1990), meta-learning is another line of work that shares a similar spirit with learning to optimize (Gharoun et al., 2023). A large proportion of works on meta-learning target few-shot learning tasks, which prepare the model, not the optimizer, for downstream tasks (Vinyals et al., 2016; Finn et al., 2017; Yu et al., 2024b; Sun et al., 2019). Among them, Meta-SGD (Li et al., 2017) is noteworthy, as it prepares the optimizer. However, the problem set we address is general gradient-based learning, which differs from the tasks of concern in meta-learning.